### v1

# Are AI Detectors Good Enough? A Survey on Quality of Datasets With Machine-Generated Texts

Abstract
The rapid development of autoregressive Large Language Models (LLMs) has significantly improved the quality of generated texts, necessitating reliable machine-generated text detectors. A huge number of detectors and collections with AI fragments have emerged, and several detection methods even showed recognition quality up to 99.9% according to the target metrics in such collections. However, the quality of such detectors tends to drop dramatically in the wild, posing a question: Are detectors actually highly trustworthy or do their high benchmark scores come from the poor quality of evaluation datasets? In this paper, we emphasise the need for robust and qualitative methods for evaluating generated data to be secure against bias and low generalising ability of future model. We present a systematic review of datasets from competitions dedicated to AI-generated content detection and propose methods for evaluating the quality of datasets containing AI-generated fragments. In addition, we discuss the possibility of using high-quality generated data to achieve two goals: improving the training of detection models and improving the training datasets themselves. Our contribution aims to facilitate a better understanding of the dynamics between human and machine text, which will ultimately support the integrity of information in an increasingly automated world.
Are AI Detectors Good Enough? A Survey on Quality of Datasets With Machine-Generated Texts
German Gritsaithanks: Equal contributions Advacheck OÜ, Tallinn, Estonia Université Grenoble Alpes, Grenoble, France gritsai@advacheck.com                        Anastasia Voznyuk Advacheck OÜ, Tallinn, Estonia voznyuk@advacheck.com
Andrey Grabovoy Advacheck OÜ, Tallinn, Estonia grabovoy@advacheck.com                        Yury Chekhovich Advacheck OÜ, Tallinn, Estonia chekhovich@advacheck.com
1 Introduction
The quality of the outputs of autoregressive Large Language Models (LLM) has grown tremendously in the last five years, making their output almost indistinguishable from human-written texts Chang et al. (2024). This expanded the application fields of such models, as many routine tasks can be entrusted to them nowadays. However, we may observe multiple cases of overuse of these models, when they are utilised for creating texts that are intended to be written and fact-checked by humans. Misuse could be revealed in the generation of fake news Zellers et al. (2019); Zhou et al. (2023), which can mislead readers of such content. Another example is concern in academic society, where a lot of students complete assignments with LLMs Koike et al. (2024); Ma et al. (2023), making professors grade AI-generated content and undervaluing the purpose of the educational process. The spread of machine-generated fragments and plagiarism in scholarly articles increases rapidly with the growth of chatbots and reaches several tens of percent Liang et al. (2024); Gritsay et al. (2023a). Field researches also showed that in the last year alone, more than 60,000 scientific papers contain evidence of the use of machine generation Gray (2024). As generated texts are gaining rapid and widespread popularity, it is crucial to develop systems able to counter the uncontrolled proliferation of artificial data and signal to the reader that the content they read is generated.
Another concern is that the Web is overflowing with machine-generated content, often of poor quality. The fact that such texts contribute bias to publicly available texts on the Internet, through false facts, hallucinations and spelling errors, is worth considering. Given the current agenda of using trillions of tokens from the Internet to teach new language models, Villalobos et al. (2022) revealed that the data will run out by 2028. That means that the training sets for more advanced language models in the future will include a large amount of generated content, as it already has a place in the current picture of available texts from all over the Internet. It was shown in Alemohammad et al. (2023) that such self-consuming will result in substantial degradation of the model’s abilities. Furthermore, the trend is evolving in such a way that human-written texts on almost any topic will be much harder to retrieve. While for texts dated 10 years ago we are confident that the usage of generation was extremely rare, we cannot state the same for modern texts.
Detectors capable to distinguish human-written texts from AI-generated texts, and whose detection quality can be guaranteed, are required by numerous specialists in many fields. We consider that the key to success in building excellent detectors will be high-quality artificial text collections. In this paper we would like to estimate the quality of the available generated data from competitions and research papers. As we are still in a lack of qualitative detectors, even though we see the claimed metrics on various datasets up to 99.9%, meanwhile in the wild a noticeable decline in performance can be seen. Competitions are a rather good way to suggest attractive challenges for researchers, but the results, obtained in these competitions may look confusing: as the models become more and more advanced, seemingly making the detecting task more challenging, participants of competitions still reach almost perfect scores of metric, bringing up the problem about quality of generated data in the provided datasets, see Appendix  C.
Our contributions are following:
  1. 1.
    We systemize information about datasets from articles and competitions, dedicated to the detection of AI-generated content task.
  2. 2.
    We suggest methods that may be helpful for evaluating the quality of the generated data and the datasets aimed to use for binary classification between human and machine texts.
2 Related Work
AI-generated content detection is usually a text classification task, meaning that the input is a text sequence and the output is a discrete, usually binary class prediction. When the task is binary, the common labels are “AI” or “human”, however in some cases one may be interested in multi-class classification between several language models. The last task is usually called authorship attribution. Finally, there is a recent, more complex task, called hybrid writing detection, when text contains fragments of both authors and the task is to determine the borders when authors change.
Among the approaches to tackle the classification problem Jawahar et al. (2020) is to calculate linguistic Fröhling and Zubiaga (2021), stylometric, and statistical features, as well as using classical machine learning methods such as Logistic Regression or Random Forest as classifiers on these features. Some other techniques include the calculation of internal metrics, such as token-wise log probability or rank Gehrmann et al. (2019). They can be calculated for each token and compared for consistency with the prior context. However, these methods are falling behind approaches that use BERT-based models and keep an eye on the context Ippolito et al. (2020); Gritsay et al. (2022) such as RoBERTa Liu et al. (2019) and DeBERTa He et al. (2021a), that are showing state-of-the-art results on this task Uchendu et al. (2021).
Regarding methods for evaluating the quality of the generated data itself, it has become more common to evaluate with the help of LLMs themselves Xu et al. (2023) . The main advantage of such approach is that it does not require any human reference, unlike ROUGE Lin (2004) or MoverScore Zhao et al. (2019). However, the output of model-evaluator needs to be unified and is not always interpretable. Another approach is suggested by Zhu and Bhat (2020), where the text is evaluated based on several linguistic criteria, such as grammar or coherence.
There are many datasets, developed as benchmarks for AI-generated content detection and it would be infeasible to evaluate them all. We will describe the datasets we used in our analysis and experiments in Section 3. There are a number of works dedicated to summarization of information about those datasets, however, only few works focus on the quality of data in them. Participants of competitions often provide a brief overview of the data in their solution descriptions Shamardina et al. (2022b); Rosati (2022); Gritsay et al. (2023b); Boeva et al. (2024); Zhao et al. (2024); Gritsai et al. (2024); Miralles et al. (2024); Valdez-Valenzuela et al. (2024). In Voznyuk and Konovalov (2024) it was shown that in SemEval Task 8 Wang et al. (2024b) there were a number of generations of poor quality, which made the task for detectors much easier. In this work, we analyse some other datasets, whether they also contain similar sort of generation errors, by diminishing the quality of detectors trained on them.
3 Data
3.1 Datasets From Shared Tasks
Shared tasks to detect artificially generated fragments appeared several years ago. Participants are suggested to solve some tasks connected with detection, either a binary classification, authorship attribution or hybrid human-AI collaboration. In this work, we focused on competitions in which a binary classification task was present. It should be noted that the studies only present data that were made available for training stage, as not all competitions had gold labels provided for test samples after the end of the competition. All shared tasks contain text in English, unless stated otherwise:
RuATD 2022 Shamardina et al. (2022a) contains human- and machine-written fragments in Russian with wide range of themes.
DAGPap 2022 Kashnitsky et al. (2022) contains human- and machine-written scientific excerpts collected by Elsevier.
AuTexTification 2023 Sarvazyan et al. (2023) provides texts for classification in English and Spanish. The organisers considered five different domains to solve with.
IberAuTexTiification 2024 Sarvazyan et al. (2024) is a continuation of the previous competition. The novelty of this shared task is that it is multilingual (six Iberian languages), multi-domain and multi-model.
Voight-Kampff Generative AI Authorship Verification 2024 Ayele et al. (2024) provides two texts, one authored by a human, one by a machine, and participants were suggested to pick the human one. Test data for this task was compiled from the submissions of another competition’ participants and comprised multiple text genres. For shorter term in future mentions, we will refer to this competition as PAN 2024.
SemEval 2024 Task 8 Wang et al. (2024b) focuses on shifts between domains, generators and languages of generated texts. Among available languages are Chinese, Urdu, Bulgarian, Indonesian, Russian, Arabic, German and Italian, however test set consisted only of texts in English, Italian, German and Arabic.
MGT Detection Task 1  [^1] COLING-25 MGT Detection Task 1 is the continuation and improvement of the SemEval Task 8. Organisers aim to refresh training and testing data with generations from novel LLMs and include new languages.
A basic overview of the analysed datasets is presented in Table 1, and a more detailed description of the data source and the topics presented can be seen in Appendix B.
Dataset Year Language Num. of Texts Generated VS Human Average Length Median Length
RuATD 2022 ru 129k 64.5k / 64.5k 236.86 / 221.47 99.0 / 95.0
DAGPap22 2022 en 5.3k 3.6k / 1.6k 799.45 / 1180.07 680.0 / 1126.5
AuTex 2023 en, es 65.9k 33.1k / 32.8k 315.08 / 297.28 386.0 / 351.0
IberAuTex 2024 es, en, ca, gl, eu, pt 98k 52.5k / 45.4k 1036.92 / 1058.36 981.0 / 1018.0
PAN24 2024 en 15.2k 14.1k / 1.1k 2640.50 / 3007.04 2731.0 / 2868.0
SemEval24 Mono 2024 en 34.2k 18k / 16.2k 2465.12 / 2358.05 2570.0 / 2083.5
SemEval24 Multi 2024 en, ar, de, it 42.3k 22.1k / 20.2k 2217.87 / 2256.67 2270.0 / 2032.0
MGT Task 1 Mono 2025 en 610.7k 381.8k / 228.9k 1448.28 / 1541.18 1208.0 / 1080.0
MGT Task 1 Multi 2025 en, zh, it, ar, de, ru, bg, ur, id 674k 416.1k / 257.9k 1422.74 / 1445.33 1195.0 / 1032.0
Table 1: Statistics of the datasets from the shared tasks.
3.2 Datasets from Papers
The number of collections with generated content has started to increase with an increasing number of available generators. Quite often, researchers together with a new approach for AI content detection publish a parallel dataset on which they have validated their method. In this paper, we picked collections with human- and machine-generated excerpts that are the most common and cited in other researchers’ publications.
GPT2 Output Dataset 222gpt2-output-dataset consists of text outputs generated by GPT-2 models across various prompts.
HC3 (Human Chatbot Conversations Corpus) Su et al. (2024) features conversations between humans and chatbots, primarily used for research on chatbot responses and human-AI interaction analysis. This dataset is available for both English and Chinese, but we have focused only on the former.
GhostBuster Verma et al. (2024) aimed at detecting AI-generated content by comparing it to human-written text, often used in the context of identifying machine-generated misinformation or spam.
MGTBench (Machine Generated Text Benchmark) He et al. (2023) is a benchmark dataset designed to evaluate the quality of machine-generated text across various tasks, including fluency, coherence, and creativity.
MAGE (Model Augmented Generative Evaluation) Li et al. (2024) evaluates the performance of generative models by comparing outputs with human annotations, aiding in the development of more accurate generative AI models.
M4 (Multilingual, Multimodal, Multitask, Massive Dataset) Wang et al. (2023) is a large-scale dataset designed for training models that can handle multiple languages, tasks, and modalities, making it useful for developing versatile AI systems. It contains texts in Arabic, Bulgarian, Indonesian, Russian, Chinese and Urdu, but we focused only on texts in English.
Dataset Year Language Num. of Texts Generated VS Human Average Length Median Length
GPT2 2019 en 1250k 1000k / 250k 2941.28 / 2616.04 3245.0 / 2459.0
HC3 2023 en 85.4k 26.9k / 58.5k 1010.50 / 680.68 1012.0 / 422.0
GhostBuster 2023 en 21k 18k / 3k 3345.07 / 3391.26 3439.5 / 2911.5
MGTBench 2024 en 23.7k 20.7k / 3k 1595.94 / 3391.26 1226.0 / 2911.5
MAGE 2024 en 436k 152.3k / 284.2k 1138.75 / 1281.88 706.0 / 666.0
M4 2024 en 89.5k 44.7k / 44.7k 1587.62 / 3162.40 1454.0 / 1697.0
Table 2: Statistics of the datasets from the research papers.
4 Approach
We decided to evaluate all datasets on common setups to see how good standard approaches perform on them. We did not have the goal to obtain the highest score, but rather to compare the performance of the same method on different datasets.
4.1 Classifiers
DeBERTa Classifier. A standard approach for binary classification of documents is to fine-tune in supervised manner some classifier, usually BERT-like model. In our case, we used mDeBERTa He et al. (2021b), which is the current state-of-the-art model for machine-generated text detection Macko et al. (2023).
DetectGPT. The DetectGPT framework Mitchell et al. (2023) introduced novel perplexity-based scoring function, which involves perturbing a text passage, after what log-probabilities between original and modified texts are compared. This score is passed as an input to a classification model, with GPT-2 Radford et al. (2019) as the base model and T5-Large Raffel et al. (2019) as perturbations generator. However, as DetectGPT requires intensive computational costs, we utilized Fast-DetectGPT Bao et al. (2024), that substitutes DetectGPT’s perturbation step with a more efficient sampling step. Another advantage is it can be requested via API and approach is applicable without additional fine-tuning stages.
Binoculars Hans et al. (2024) introduce modified score with perplexities and cross-perplexity of two close language models to separate human and machine texts and is proved to be better than approach with classic perplexity. Also, the approach is applicable without additional fine-tuning stages.
4.2 Topological Time Series calculation
It was shown in Tulchinskii et al. (2023) that if we take the inner dimensionality of the manifold on the set of embeddings, we could separate human-written texts from machine-generated ones. However, in that work authors estimated short texts, and therefore we followed the approach in Kushnareva et al. (2024) with sliding window. Topological Time Series calculate intrinsic dimensions (PHD) of the text within sliding window and they can be used as a feature for detector. Authors conclude that this metric can, in fact, help to differentiate the texts of different origin. To be able to compare datasets between each other, we came up with a score, utilizing KL-divergence. Let $h_{d}$, $m_{d}$ be distributions of intrinsic dimensions for two types of texts from the same dataset, of human and machine origin, then our $\text{KL}_{\text{TTS}}$ is following:
$\text{KL}_{\text{TTS}}(h_{d},m_{d})=|D_{\text{KL}}(h_{d}||m_{d})-D_{\text{KL}}(m_{d}||h_{d})|$
The lower this score, the closer $h_{d}$ and $m_{d}$ are, which means almost indistinguishable texts and vice versa.
4.3 Perturbations and Shuffling
Based on the results of text modification studies Sadasivan et al. (2024); Mitchell et al. (2023), which show how small perturbations affect machine reading comprehension systems, we decided to consider this way of possibly assessing the quality of a dataset. The key idea here is that AI models are sensitive to such adversarial changes, unlike humans. We considered two modification ideas: Adversarial Token Perturbation and Sentence Shuffling.
Adversarial Token Perturbation. In this approach we divide the text into tokens and randomly replace the token with a synonym from the WordNet Miller (1994) collection with a probability of 70%. We apply such a technique to each represented class. Using an encoder model, we obtain embeddings for each of the texts in the current dataset. Finally, we measure the average embedding shifts for the classes of human and generated texts. We obtain the embedding shifts using the cosine distance between the embeddings of the original texts and the modified ones. As a result, after modifications we obtain $\Delta_{\text{shift}}$ — the log difference of the average embedding shifts.
$\Delta_{\text{shift}}=\log\frac{{\frac{1}{n}\sum_{i=1}^{n}\text{cos}_{d}(h_{h_{i}}^{o},h_{h_{i}}^{p})}}{{\frac{1}{m}\sum_{j=1}^{m}\text{cos}_{d}(h_{m_{j}}^{o},h_{m_{j}}^{p})}},$
where $n$ and $m$ – number of samples in the human and generated parts of the dataset respectively, $h_{h_{i}}^{o}$ – embedding of $i$-th fragment of human part of data, $h_{h_{i}}^{p}$ – the same embedding after perturbation. The same goes for $h_{m_{i}}^{o}$ and $h_{m_{i}}^{o}$ for machine-generated part of data. Finally, $\text{cos}_{d}$ is a function that measures the cosine distance between two vectors.
Sentence Shuffling. In this approach, we randomly swap sentences, thereby affecting the cohesion of the text. We try to find out the effect of artificial origin on the difference between the distributions after permutations. By dividing a fragment into sentences and randomly reversing the order of 70% of the selected sentences, we apply this technique to each represented class. Further, using the text encoding model, we obtain embeddings for each of the texts of the current dataset. Finally, we measure embedding shifts for the class of human and generated texts, and after that we convert the shifts into probability-like distributions. This allows us to obtain at the end $\text{KL}_{\text{shuffle}}(H,M)$ — the KL-divergence between the shifts of human and generated texts.
$\text{KL}_{\text{shuffle}}(H,M)=\sum_{i}H(i)\log\frac{H(i)}{M(i)},$
$H(i)=\frac{\text{cos}_{d}(h_{h_{i}}^{o},h_{h_{i}}^{p})+\epsilon}{\sum_{j}\left(\text{cos}_{d}(h_{h_{j}}^{o},h_{h_{j}}^{p})+\epsilon\right)},$
and $M(i)$ has the same structure, except that instead of human class texts the generated class texts are used, $\epsilon$ is a small constant added to avoid division by zero.
Figure 1: Comparisons of embedding shifts after two types of modifications for the HC3 dataset.
4.4 Attention Maps
We hypothesised that some patterns in the attention map might be useful in evaluating the quality of the texts, as in the case of perfect quality of generation, the performance of the model on human and machine texts should be identical. As it is infeasible to evaluate attention maps visually for large amounts of texts, we came up with a statistic that may be useful in grasping some patterns. It was intended to grasp “attention columns” or massive activations Sun et al. (2024) that appear on the attention map. For some texts there is no such pattern observed, for others there’s only one distinct column, then there are cases when two or three distinct columns are seen. Our hypothesis was that attention maps for machine-generated texts should show more columns than for human ones, because there may be some tokens that model does not expect to see.
$Attention_{j}(t)=\frac{1}{|t_{j:|t|}|}\sum_{i=j}^{|t|}Attention(q_{i},k_{j})$
Therefore, until the end of the text, it still attends to such tokens, and we calculated the mean difference between the top-1 and the top-2 values across the dataset. Depending on the number of “attention columns”, this difference will vary, but in general we assume that it should be comparable for texts of both origin.
5 Experiments
Dataset DeBERTa Binoculars DetectGPT
GPT-2 0.972 0.495 0.412
HC3 0.998 0.931 0.972
GhostBuster 0.910 0.683 0.711
MGTBench 0.961 0.164 0.344
MAGE 0.835 0.632 0.654
M4 0.987 0.171 0.381
SemEval24 Mono 0.999 0.943 0.983
SemEval24 Multi 0.997
RuATD 0.765
DAGPap22 0.968 0.333 0.562
PAN24 0.826 0.411 0.890
AuTex23en 0.941 0.783 0.911
AuTex23es 0.933
IberAuTex 0.964
MGT-1 Mono 0.904 0.665 0.683
MGT-1 Multi 0.934
Table 3: Classification results with different detectors estimated using $F_{1}$-score. Binoculars and DetectGPT work only with English texts, thus we could not apply them to datasets with non-English texts.
From each dataset, we sampled 1000 documents, balanced between two classes. As we dealt with multilingual datasets, we chose to finetune a multilingual model supporting all the presented languages, namely mDeBERTa-v3-base. Information about training can be found in the Appendix D.To evaluate the quality of baselines, Binoculars and Fast-DetectGPT, we launched falcon-rw-1b Almazrouei et al. (2023) and gpt-neo-2.7B Black et al. (2021) respectively. It is worth noting that with the last two methods we were only able to measure quality for samples in English. Our objective was to show that datasets of lower quality have shifts that will be easily recognised by the models ”from the first step”, hence we have not performed any hyperparameter tuning, only one iteration of fine-tuning and testing of the underlying models. In the experiment with topological features we used RoBERTa-base, just as the authors of original paper. In the experiment Perturbations and Shuffling, the multilingual-e5-large encoder was used to build embeddings of texts, which shows high metrics on encoding high-resource languages Wang et al. (2024a). Finally, in the experiments with attention maps we used LLaMA2-7B to obtain attention scores. We took 3 heads from Layer 15 and averaged them, as heads on this layer showed very distinct attention patterns.
6 Results
Dataset $\text{KL}_{\text{TTS}}$downarrow Attention Columns (h / m) $\Delta_{\text{shift}}\downarrow$ $\text{KL}_{\text{shuffle}}\downarrow$
GPT-2 0.014 3.430 / 4.094 0.084 1.255
HC3 0.053 0.459 / 0.967 0.264 1.167
GhostBuster 0.053 2.822 / 2.988 0.024 0.359
MGTBench 0.043 1.961 / 2.639 0.031 0.421
MAGE 0.011 2.289 / 2.166 0.094 0.310
M4 0.036 3.842 / 2.256 0.107 0.483
SemEval24 Mono 0.012 1.540 / 0.766 0.191 2.576
SemEval24 Multi 0.001 2.123 / 0.830 0.059 2.046
RuATD 0.007 1.631 / 1.391 0.315 14.028
DAGPap22 0.083 0.637 / 0.675 0.039 0.472
PAN24 0.053 3.463 / 2.588 0.050 0.331
AuTex23-en 0.021 3.179 / 2.740 0.110 4.331
AuTex23-es 0.001 3.072 / 3.244 0.105 1.306
IberAuTex 0.012 2.049 / 1.946 0.223 5.516
MGT-1 Mono 0.019 2.070 / 1.783 0.031 0.587
MGT-1 Multi 0.006 3.313 / 3.117 0.027 0.522
Table 4: Calculated statistics on texts from chosen datasets. In “Attention Columns” we show the mean difference between the highest attention column and the second-placed. The first value is for human texts; the second value is for machine-generated. Some values for $\text{KL}_{\text{TTS}}$ are underlined, because texts are too short, see Section 7.
Figure 2: Topological Time Series for different datasets. These datasets obtained highest $\text{KL}_{\text{TTS}}$. The results for the remaining datasets selected in this paper can be found in Figure 4.
Figure 3: Attention maps on two excerpts from DAGPap22, Layer 15, Head 15.
The results of comparison of the designed features on the test parts of the selected datasets are presented in Table 4. Regarding the TTS score, previous works have shown that texts of different origin have different PHD values, however this result was obtained for GPT-2 model, meanwhile current advanced models generate texts of much better quality, therefore PHD values become more similar. If texts of different origin have high $\text{KL}_{\text{TTS}}$, it means that it is easier for a detector to separate such texts. $\text{KL}_{\text{TTS}}$ is also constrained for shorter texts, see Section 7.
The second column of Table 4 is based on attention map analyses. Here, on the left are the mean difference values of the human excerpts, on the right are the generated ones. When the attention maps of texts of different origin show different patterns, it also becomes easier for detectors to separate the texts. In the remaining columns we list the statistics observed on modified texts, and for both of these the lower the better, as this reflects the similar degree of resilience of the generated and human texts to adversarial attacks. Qualitatively generated data with no bias should take values close to human.
Finally, in Table 3 we show the results of applying modern classifiers to the chosen test parts. For instance, on the datasets that had low values in Table 4, a quality close to 1 can be achieved, which indicates the clear presence of detector bias towards them, or a structural feature that is too obvious for the model. It is not possible to judge the quality of the data only by achieving $F_{1}$ values close to 1, but by combining the values of the two tables we can estimate which set has better quality data and which has lower quality data.
7 Discussion
Regarding $\text{KL}_{\text{TTS}}$, on Fig. 2 we show 4 datasets with high value of it. While GhostBuster and PAN24 received such high score due to discrepancy on texts with higher dimensions, MGTBench and DAGPap22 did it due to the difference in distributions themselves. Note also that $\text{KL}_{\text{TTS}}$ may not perform well with very short texts, since the internal method of computing PHD requires sufficiently long texts for stable computation. Therefore, we discard $\text{KL}_{\text{TTS}}$ on RuATD and AuTex23-es, as they do not fit the criteria, see Table 1. On top of that, it has already been shown that texts must be of sufficient length Gritsay et al. (2022) to build reliable detectors.
Regarding attention maps, in Fig. 3, we depict an example of distinct patterns where human text has only one separate attention column besides attention on the first token, while machine-generated text has multiple attention columns.
Analysing the values in the table 4, we can trace the presence of sufficiently high quality data in the selected datasets. The developed attributes in aggregate are able to reflect the quality of the generated dataset from different perspectives and angles. We propose to utilise these attributes in combination with other statistical tools for evaluating data quality, e.g. Zipf’s law Powers (1998).
Presented statistics can be utilised to estimate the quality of collections and to improve them. Also, datasets that collect machine-generated content may provide utility for the more general two other purposes as well. First, high-quality generated data can be utilised to evaluate the quality of the causal model during training, as one of the training objectives to improve model answers and make it more human-like. Secondly, good detectors can help to clean training sets, as large proportion of low-quality generated texts in those sets can result in emerging biases towards incorrect structure and rubbish fragments in the output of the model in the future.
8 Conclusion
In the current research, we discussed the problem of quality of datasets with AI-generated texts used for testing corresponding detectors. This problem is relevant, as the quality of test data directly influences the quality of widely used detectors. We conducted a review of datasets from competitions and scientific publications on datasets aimed at the detection of AI-generated content and proposed methods to evaluate the quality of datasets containing AI excerpts based on different structural features. We evaluated topological features, robustness to adversarial attacks, and similarity of attention patterns as estimators of data quality. We concluded that all analysed datasets fail in one another of our methods and do not allow to reliably estimate AI detectors. We encourage researchers to propose their own ways for quality assessment, which will allow to create a comprehensive system of evaluation of the detection datasets. Our work aims to contribute to a better understanding of the difference between human and machine text, which will ultimately contribute to preserving the integrity of information in the world.
9 Limitations
In our work we focused on the task of binary classification, thus suggested methods are not optimal for the task of detection of the hybrid AI-human content. Also, some methods do not work properly on short texts, however, so do the detectors.
Appendix A Additional Results on Topological Time Series
In Figure 4, it can be seen the plotted Topological Time Series distributions for all the selected datasets for the currect research.
Figure 4: Topological Time Series for remaining datasets from current review. We omitted the results for AuTex23en, because virtually all texts there had the dimension of 0.
Appendix B Data Description
More detailed description with information on sources, topics and years of the datasets selected in this paper from competitions in Table 5 and research articles presented in Table  6.
Dataset Year Themes Sources
GPT2 2019 WebText GPT-2-117M, GPT-2-345M, GPT-2-762M, GPT-2-1542M
HC3 2023 ELI5, WikiQA, Wikipedia, Medicine, Finance ChatGPT (gpt-turbo-3.5)
GhostBuster 2023 Student Essays, News Articles, Creative Writing ChatGPT (gpt-3.5-turbo), Claude
MGTBench 2024 Student Essays, News Articles, Creative Writing ChatGLM, Dolly, ChatGPT-turbo, GPT4All, StableLM, Claude
MAGE 2024 Opinions, Reviews, News, QA, Story Generation, Commonsense Reasoning, Knowledge Illustration, Scientific Writing ChatGPT (text-davinci-002, text-davinci-003, gpt-turbo-3.5), LLaMA, GLM-130B, FLAN-T5, OPT, BigScience, EleutherAI
M4 2024 Wikipedia, Reddit ELI5, WikiHow, PeerRead, arXiv abstract ChatGPT (text-davinci-003, gpt-turbo-3.5), Cohere, Dolly-v2, BLOOMz
Table 6: More detailed descriptive statistics about domains and generators of the chosen datasets from papers.
Appendix C Evaluation results of competitions
Table 7 shows the winning scorers in the competitions reviewed in this paper. In the AuTex and IberAuTex competitions it was forbidden to use additional data to fine-tune the detection algorithms. In the other collections it was allowed, we can notice a high quality near perfect in them. We should note the low value of metrics on the RuATD dataset, which can be explained by the limited number of available high-quality language models in Russian during the competition.
Competition Metric Best result
RuATD Accuracy 0.820
AuTex-en Macro-F1 0.809
AuTex-es Macro-F1 0.708
IberAuTex Macro-F1 0.805
SemEval24 Mono Accuracy 0.975
SemEval24 Multi Accuracy 0.959
PAN24 Avg. of 5 metrics* 0.924
DAGPap22 Avg. F1-score 0.994
Table 7: Best results from each analysed competition. PAN24 used mean of 5 metrics, such as accuracy, F1 and other to evaluate efficiency of the system.
Appendix D Hyperparameters
Hyperparameters Values
Epochs 5*
Learning rate (LR) 5e-5
Warmup steps 50
Weight decay 0.01
Table 8: Hyperparameters for fine-tuning mDeBERTa-base. We trained for 5 epochs with possibility of early exit.
The training was conducted on NVIDIA GeForce RTX 3090. See hyperparameters in Table 8.