### v2

# Are AI Detectors Good Enough? A Survey on Quality of Datasets With Machine-Generated Texts

Abstract
The rapid development of autoregressive Large Language Models (LLMs) has significantly improved the quality of generated texts, necessitating reliable machine-generated text detectors. A huge number of detectors and collections with AI fragments have emerged, and several detection methods even showed recognition quality up to 99.9% according to the target metrics in such collections. However, the quality of such detectors tends to drop dramatically in the wild, posing a question: Are detectors actually highly trustworthy or do their high benchmark scores come from the poor quality of evaluation datasets? In this paper, we emphasise the need for robust and qualitative methods for evaluating generated data to be secure against bias and low generalising ability of future model. We present a systematic review of datasets from competitions dedicated to AI-generated content detection and propose methods for evaluating the quality of datasets containing AI-generated fragments. In addition, we discuss the possibility of using high-quality generated data to achieve two goals: improving the training of detection models and improving the training datasets themselves. Our contribution aims to facilitate a better understanding of the dynamics between human and machine text, which will ultimately support the integrity of information in an increasingly automated world.
Are AI Detectors Good Enough? A Survey on Quality of Datasets With Machine-Generated Texts
German Gritsaithanks: Equal contributions Advacheck OÜ, Tallinn, Estonia Université Grenoble Alpes, Grenoble, France gritsai@advacheck.com                        Anastasia Voznyuk Advacheck OÜ, Tallinn, Estonia voznyuk@advacheck.com
Andrey Grabovoy Advacheck OÜ, Tallinn, Estonia grabovoy@advacheck.com                        Yury Chekhovich Advacheck OÜ, Tallinn, Estonia chekhovich@advacheck.com
1 Introduction
The quality of large language models (LLMs) has grown tremendously in the last five years, making their output almost indistinguishable from human-written texts  Chang et al. (2024). This expanded the application fields of these models, as many routine tasks can be entrusted to them nowadays. However, they can be used for creating texts that are intended to be written and fact-checked by humans. Example of such misuse is generation of fake news Zellers et al. (2019); Zhou et al. (2023), which can mislead readers of generated content. Teachers raise another concern, as many students complete assignments with LLMs Koike et al. (2024); Ma et al. (2023), undervaluing the purpose of the educational process. Machine-generated fragments also appear in academic articles more often with the growth of chatbots and reach several tens of percent  Liang et al. (2024); Gritsay et al. (2023a). More than 60,000 scientific papers in the last year alone contained the evidence of the use of machine generation Gray (2024). All of that prove that it is crucial to develop systems able to counter the misuse of artificial data and signal to the reader that the content they read is generated.
Another concern is the Web, overflowing with machine-generated content, often of poor quality. Such texts contribute bias to publicly available texts on the Internet, through false facts, hallucinations and spelling errors. Given the current agenda of using texts from the Internet to train new language models, all this bias will be inadvertently added to the model. Moreover,  Villalobos et al. (2022) revealed that the human-written data will run out by 2028. That means that the training sets for language models in the future will include a large amount of generated content. Such self-consuming will result in the substantial degradation of the model’s abilities Alemohammad et al. (2023). Furthermore, the trend is evolving in such a way that human-written texts on almost any topic will be much harder to retrieve. While for texts dated even 5 years ago we are confident as the usage of generation was extremely rare, we cannot state the same for more recent texts.
Therefore, detectors, capable to distinguish human-written texts from AI-generated texts, and whose detection quality can be guaranteed, are necessary for many fields. We believe one of the key factors for building reliable detectors is the high-quality artificial text collections that can be used for training and evaluation. In this paper, we would like to estimate the quality of the available generated texts from competitions and research papers. Sometimes we see that some methods from participants of the competitions reach almost perfect (up to 99.9%) metric score, meanwhile in the wild we observe a noticeable decline in performance. Such results look confusing, because the models become more and more advanced, seemingly making the detecting task more challenging, meanwhile participants of competitions still reach almost perfect scores, bringing up the question about quality of generated data in the provided datasets. Are the devised methods really good or is the data easy enough for detectors to solve the seemingly hard detection task?
Our contributions are as follows.
  1. 1.
    We systemize information about existing datasets from the research papers and competitions, dedicated to the detection of AI-generated content task.
  2. 2.
    We suggest methods that may be helpful for evaluating the quality of the generated data and the datasets aimed to use for binary classification between human and machine texts.
2 Related Work
2.1 The Task of AI-generated text detection
The task of AI-generated text detection task is generally stated as a text classification task, which means that the input is a text sequence and the output is a discrete, usually binary, class prediction. When the task is binary, the common labels are “AI” or “human”, whereas multiclass classification focusses on distinguishing several language models. The last task is usually called authorship attribution. Finally, more complex task suggests to determine the borders betweens fragments from different authors, for example between human author and some LLM author.
The approach to tackle the classification problem is to utilize linguistic, stylometric, and statistical features for classifiers Jawahar et al. (2020); Fröhling and Zubiaga (2021). Some internal metrics, such as perplexity or its modifications Hans et al. (2024) can also be helpful as a inference method. Also, perturbing texts can also provide valuable information, as described in DetectGPT Mitchell et al. (2023) method, where log-probabilities between between original and modified texts are compared. Finally, finetuning encoder-based model, like DeBERTa He et al. (2021), is considered the best solution for this task Uchendu et al. (2021); Macko et al. (2023).
2.2 Evaluating Generated Text
As for evaluating the quality of the generated data itself, it has become more common to evaluate it with the help of LLMs Xu et al. (2023). This approach does not require any human reference, unlike ROUGE Lin (2004). However, the downside is the output of model-evaluator needs to be unified, not always interpretable and model-evaluator scores can be skewed. Alternative approach is suggested by  Zhu and Bhat (2020), where the text is evaluated based on several linguistic criteria, such as grammar or coherence.
2.3 Datasets with Artificial Content
There are a number of surveys about machine-generated content detection with an overview of the datasets Jawahar et al. (2020); Wu et al. (2023), however, few works focus on the quality of data in the available datasets, despite it being an important aspect of the task. Building AI-generated content detectors requires high-quality labelled data that involve substantial financial, computational, and human resources. The human evaluators should check that the dataset does not contain corrupted generations, that the texts are coherent and grammatically correct. We will describe the datasets we used in our analysis and experiments in Section 3.
2.4 Shared Tasks on AI-Generated content Detection
Shared tasks help to advance research on detecting AI-generated content, offering new variations on the tasks and data for evaluation, encouraging participants to come up with novel ideas for detectors robust to the change of language, domain, or generating model. Participants explore approaches ranging from transfer learning on complex text features to utilising and fine-tuning LLMs for these tasks (2022; 2022; 2023b; 2024; 2024; 2024; 2024; 2024). These efforts have highlighted challenges such as handling multilingual data and adapting to rapidly evolving generative models. Some participants also provide some analysis of the given data or even discuss some flaws with generated texts in it Voznyuk and Konovalov (2024).
3 Datasets
Dataset Year Language # Texts, $10^{3}$ Texts, G / H, $10^{3}$ Average Length, G / H Median Length, G / H
GPT2 2019 en 1250 1000 / 250 2941 / 2616 3245 / 2459
TweepFake 2019 en 20.7 10.4 / 10.4 104 / 118 89 / 94
HC3 2023 en, zh 85.4 26.9 / 58.5 1011 / 681 1012 / 422
GhostBuster 2023 en 21 18 / 3 3345 / 3391 3440 / 2911.5
MGTBench 2024 en 23.7 20.7 / 3 1596 / 3391 1226 / 2911.5
MAGE 2024 en 436 152.3 / 284.2 1139 / 1282 706 / 666
M4 2024 en, zh, ru, bg, ur, id 89.5 44.7 / 44.7 1588 / 3162 1454 / 1697
OutFox 2024 en 57.6 43.2 / 14.4 2686 / 2238 2311 / 1992
DAGPap22 2022 en 5.3 3.6 / 1.6 799 / 1180 680 / 1126.5
RuATD 2022 ru 129 64.5 / 64.5 237 / 221 99 / 95
AuTex 2023 en, es 65.9 33.1 / 32.8 315 / 297 386 / 351
IberAuTex 2024 es, en, ca, gl, eu, pt 98 52.5 / 45.4 1037 / 1058 981 / 1018
PAN24 2024 en 15.2 14.1 / 1.1 2641 / 3007 2731 / 2868
SemEval24 Mono 2024 en 34.2 18 / 16.2 2465 / 2358 2570 / 2083.5
SemEval24 Multi 2024 en, ar, de, it 42.3 22.1 / 20.2 2218 / 2257 2270 / 2032
MGT-1 Mono 2025 en 610.7 381.8 / 228.9 1448 / 1541 1208/ 1080
MGT-1 Multi 2025 en, zh, it, ar, de, ru, bg, ur, id 674 416.1 / 257.9 1423 / 1445 1195 / 1032
Table 1: Statistics of the texts in the datasets from the shared tasks and research papers.
3.1 Datasets From Shared Tasks
The most common tasks in shared tasks are binary classification and authorship attribution, with binary classification being the prevalent task, therefore, in this work, we focused only on it. All chosen shared tasks contain texts in English, unless stated otherwise. Here we give a brief overview of each task, as well as some quantitative statistics of the texts in Table 1, whereas a more detailed description, such as models used for generation or domains of the presented texts, can be found in Appendix A.
  1. 1.
    DAGPap 2022 Kashnitsky et al. (2022) introduced a dataset of human- and machine-written scientific excerpts collected by Elsevier.
  2. 2.
    RuATD 2022 Shamardina et al. (2022) focused on human- and machine-written documents in Russian, covering a wide range of themes.
  3. 3.
    AuTexTification 2023 Sarvazyan et al. (2023) provided texts in English and Spanish, covering five distinct domains.
  4. 4.
    IberAuTexTiification 2024 Sarvazyan et al. (2024) expanded on the previous competition with a multilingual (six Iberian languages), multi-domain, and multi-model focus.
  5. 5.
    Voight-Kampff Generative AI Authorship Verification 2024 Ayele et al. (2024), hereafter referred to as PAN 2024, tasked participants with identifying the human-authored text from two samples – one human-written and one machine-generated.
  6. 6.
    SemEval 2024 Task 8 Wang et al. (2024c) addressed domain, generator, and language shifts in generated texts. Training data included multiple languages (e.g., Chinese, Urdu, and Russian), but the test set was limited to English, Italian, German, and Arabic.
  7. 7.
    MGT Detection Task 1 (COLING 2025) Wang et al. (2025) was built on SemEval 2024 Task 8 by incorporating data generated by novel LLMs and expanding multilingual coverage of train and test sets.
3.2 Datasets from Papers
The number of collections with generated content has started to grow with an increasing number of available generators. Quite often, researchers, together with a new approach for AI content detection, publish a parallel dataset on which they have validated their method. In this paper, our aim was to pick collections with human- and machine-generated excerpts that are the most common and cited in other researchers’ publications. Similarly to previous subsection, here we give a brief overview of each chosen datasets, describe some statistics about the texts in Table 1, and add a more detailed description in Appendix A.
GPT2 Output Dataset [^1] https://github.com/openai/gpt-2-output-dataset consists of text outputs generated by GPT-2 models of different sizes across various prompts.
HC3 (Human Chatbot Conversations Corpus) Su et al. (2024) features conversations between humans and chatbots, primarily used for research on chatbot responses and human-AI interaction analysis. This dataset is available for both English and Chinese, but we have focused only on the former.
GhostBuster Verma et al. (2024) aimed at detecting AI-generated content by comparing it to human-written text, often used in the context of identifying machine-generated misinformation or spam.
MGTBench (Machine Generated Text Benchmark) He et al. (2023) is a benchmark dataset designed to evaluate the quality of machine-generated text across various tasks, including fluency, coherence, and creativity.
MAGE (Model Augmented Generative Evaluation) Li et al. (2024) evaluates the performance of generative models by comparing outputs with human annotations, aiding in the development of more accurate generative AI models.
M4 (Multilingual, Multimodal, Multitask, Massive Dataset) Wang et al. (2024b) is a large-scale dataset designed for training models that can handle multiple languages, tasks, and modalities, making it useful for developing versatile AI systems. Although it is multilingual, we sampled only English texts.
TweepFake Fagni et al. (2021) contains real tweets written by humans and synthetic tweets, generated by various AI models, from bots, imitating human users.
Outfox Koike et al. (2024) contains triplets of essay problem statements, human-written essays, and LLM-generated essays. The students who wrote the essays range from 6th to 12th grade in the USA.
Figure 1: Comparison of embedding shifts after two types of modifications for the HC3 dataset.
Figure 2: Topological Time Series for different datasets. The results for the remaining datasets selected in this paper can be found in Figure 4.
4 Approach
We decided to evaluate all datasets with common setups to see how good standard approaches perform on them. We did not have the goal to obtain the highest score, but rather to compare the performance of the same method on different datasets.
4.1 Baselines
We used several common classifiers as baselines on each dataset to obtain first-hand understanding of the dataset. The first one is mDeBERTa He et al. (2021), which is the current state-of-the-art model for multilingual machine-generated text detection Macko et al. (2023). The next baseline is DetectGPT framework Mitchell et al. (2023). We used GPT-2 Radford et al. (2019) as the base model and T5-Large Raffel et al. (2019) as perturbations generator. However, as DetectGPT requires intensive computational costs, we utilized Fast-DetectGPT Bao et al. (2024) that substitutes DetectGPT’s perturbation step with a more efficient sampling step. Finally, we used Binoculars Hans et al. (2024) as a method with improved perplexity score. The last two baselines require no fine-tuning which is crucial for detection task, as it is infeasible to train the detector on all existing artificial texts.
4.2 Topological Statistics
It was shown in Tulchinskii et al. (2023) that if we take the inner dimensionality of the manifold on the set of embeddings, we could separate human-written texts from machine-generated ones. The authors used persistence homology dimension (PHD) and showed that statistically human-generated texts have higher PHD than machine-generated texts, therefore introducing a novel detector. We calculated PHD on each set of texts. Additionally, in  Kushnareva et al. (2024) it was suggested to calculate PHD within sliding window. These intrinsic dimensions of the text within sliding window can be used as a feature for detector. The authors conclude that this metric as well can help to differentiate texts of different origin. To be able to compare datasets between each other, we came up with a score, utilizing KL-divergence, but we also needed it to be symmetrical. Let $h_{d}$, $m_{d}$ be distributions of intrinsic dimensions for two types of texts from the same dataset, of human and machine origin, then our $\text{KL}_{\text{TTS}}$ is following:
$\text{KL}_{\text{TTS}}(h_{d},m_{d})=|D_{\text{KL}}(h_{d}||m_{d})-D_{\text{KL}}(m_{d}||h_{d})|$
The lower this score, the closer $h_{d}$ and $m_{d}$ are, which means almost indistinguishable texts and vice versa.
4.3 Perturbations and Shuffling
Based on the results of the text modification studies Sadasivan et al. (2024); Mitchell et al. (2023), which show how small perturbations affect machine reading comprehension systems, we decided to consider this way of possibly assessing the quality of a dataset. The key idea here is that AI models are sensitive to such adversarial changes, unlike humans. We considered two modification ideas: Adversarial Token Perturbation and Sentence Shuffling.
Adversarial Token Perturbation. In this approach we divide the text into tokens and randomly replace the token with a synonym from the WordNet collection Miller (1994) with a probability of 70%. We apply such a technique to each represented class. Using an encoder model, we obtain embeddings for each of the texts in the current dataset. Finally, we measure the average embedding shifts for the classes of human and generated texts. We obtain the embedding shifts using the cosine distance between the embeddings of the original texts and the modified ones. As a result, after modifications we obtain $\Delta_{\text{shift}}$ — the log difference of the average embedding shifts.
$\Delta_{\text{shift}}=\log\frac{{\frac{1}{n}\sum_{i=1}^{n}\text{cos}_{d}(h_{h_{i}}^{o},h_{h_{i}}^{p})}}{{\frac{1}{m}\sum_{j=1}^{m}\text{cos}_{d}(h_{m_{j}}^{o},h_{m_{j}}^{p})}},$
where $n$ and $m$ – number of samples in the human and generated parts of the dataset respectively, $h_{h_{i}}^{o}$ – embedding of $i$-th fragment of human part of data, $h_{h_{i}}^{p}$ – the same embedding after perturbation. The same goes for $h_{m_{i}}^{o}$ and $h_{m_{i}}^{o}$ for machine-generated part of data. Finally, $\text{cos}_{d}$ is a function that measures the cosine distance between two vectors.
Sentence Shuffling. In this approach, we randomly swap sentences, thereby affecting the cohesion of the text. We try to find out the effect of artificial origin on the difference between the distributions after permutations. By dividing a fragment into sentences and randomly reversing the order of 70% of the selected sentences, we apply this technique to each represented class. Further, using the text encoding model, we obtain embeddings for each of the texts of the current dataset. Finally, we measure embedding shifts for the class of human and generated texts, and after that we convert the shifts into probability-like distributions. This allows us to obtain at the end $\text{KL}_{\text{shuffle}}(H,M)$ — the KL-divergence between the shifts of human and generated texts.
$\text{KL}_{\text{shuffle}}(H,M)=\sum_{i}H(i)\log\frac{H(i)}{M(i)},$
$H(i)=\frac{\text{cos}_{d}(h_{h_{i}}^{o},h_{h_{i}}^{p})+\epsilon}{\sum_{j}\left(\text{cos}_{d}(h_{h_{j}}^{o},h_{h_{j}}^{p})+\epsilon\right)},$
and $M(i)$ has the same structure, except that instead of human class texts the generated class texts are used, $\epsilon$ is a small constant added to avoid division by zero.
5 Experiments
From each dataset, we sampled 1000 documents from the test set, balanced between two classes. Regarding baselines, we fine-tuned mdeberta-v3-base for each dataset and evaluated the model, information about training can be found in the Appendix C. To evaluate the quality of baselines, Binoculars and Fast-DetectGPT, we utilised falcon-rw-1b Almazrouei et al. (2023) and gpt-neo-2.7B Black et al. (2021) respectively. It is worth noting that with the last two methods we were only able to measure quality for samples in English.
Our objective was to show that datasets of lower quality have shifts that will be easily recognised by the models ”from the first step”, hence we have not performed any hyperparameter tuning, only one iteration of fine-tuning and testing of the underlying models. In the experiment with topological features we used roberta-base, just as the authors of the original paper. In the experiment with perturbations and shuffling, the multilingual-e5-large encoder was used to build embeddings of texts, which shows high metrics on encoding high-resource languages Wang et al. (2024a).
6 Results
Dataset $\text{KL}_{\text{TTS}}$downarrow $\text{PHD}_{\text{human}}$ $\text{PHD}_{\text{machine}}$ $\Delta_{\text{shift}}$downarrow $\text{KL}_{\text{shuffle}}$downarrow
GPT-2 0.014 9.23 $\pm$ 1.98 10.27 $\pm$ 1.84 0.084 1.255
HC3 0.053 8.76 $\pm$ 1.83 7.38 $\pm$ 1.05 0.264 1.167
GhostBuster 0.053 9.84 $\pm$ 1.18 9.76 $\pm$ 1.15 0.024 0.359
MGTBench 0.043 8.77 $\pm$ 1.31 9.97 $\pm$ 1.02 0.031 0.421
MAGE 0.011 9.8 $\pm$ 2.14 9.38 $\pm$ 3.04 0.094 0.310
M4 0.036 7.26 $\pm$ 1.99 8.59 $\pm$ 1.4 0.107 0.483
OutFox 0.025 8.96 $\pm$ 1.21 11.48 $\pm$ 1.13 0.095 0.237
TweepFake - 9.02 $\pm$ 3.19 8.12 $\pm$ 4.02 0.116 1.001
SemEval24 Mono 0.012 9.11 $\pm$ 1.19 9.41 $\pm$ 1.2 0.191 2.576
SemEval24 Multi 0.001 9.65 $\pm$ 1.81 9.42 $\pm$ 1.44 0.059 2.046
RuATD 0.007 7.33 $\pm$ 1.4 7.46 $\pm$ 1.41 0.315 14.028
DAGPap22 0.083 8.35 $\pm$ 1.33 7.48 $\pm$ 2.01 0.039 0.472
PAN24 0.053 9.4 $\pm$ 1.05 8.52 $\pm$ 1.59 0.050 0.331
AuTex23 Eng 0.021 8.07 $\pm$ 2.26 8.1 $\pm$ 2.68 0.110 4.331
AuTex23 Esp 0.001 9.16 $\pm$ 3.49 9.25 $\pm$ 3.26 0.105 1.306
IberAuTex 0.012 9.33 $\pm$ 2.45 8.47 $\pm$ 2.73 0.223 5.516
MGT-1 Mono 0.019 9.19 $\pm$ 1.75 8.96 $\pm$ 2.24 0.031 0.587
MGT-1 Multi 0.006 8.76 $\pm$ 1.85 8.6 $\pm$ 2.29 0.027 0.522
Table 2: Calculated statistics on texts from chosen datasets. Some values for $\text{KL}_{\text{TTS}}$ are underlined, because texts are too short, see Section 7 and TTS for almost all texts in TweepFake is equal to 0.
Dataset BERT Binoculars DetectGPT
GPT-2 0.972 0.495 0.412
TweepFake 0.941 0.845 0.864
HC3 0.998 0.931 0.972
Ghost- Buster 0.910 0.683 0.711
MGT- Bench 0.961 0.364 0.447
MAGE 0.835 0.632 0.654
M4 0.987 0.871 0.881
OutFox 0.901 0.692 0.707
SemEval Mono 0.991 0.913 0.924
SemEval Multi 0.994
RuATD 0.765
DAGPap 0.968 0.333 0.562
PAN24 0.826 0.411 0.890
AuTex En 0.941 0.783 0.911
AuTex Es 0.933
IberAuTex 0.964
MGT-1 Mono 0.904 0.665 0.683
MGT-1 Multi 0.934
Table 3: Classification results with different detectors estimated using $F_{1}$-score. BERT here is DeBERTa-based classifier. Binoculars and DetectGPT work only with English texts, thus we could not apply them to datasets with non-English texts.
The results of the comparison of the designed features on the selected datasets are presented in Table 2. Regarding the PHD and TTS score, in previous works it was shown that texts from language models have smaller PHD values than human-written ones, however this result was obtained for GPT-2, GPT-3.5 and OPT models, and this trend could change for more recent language models, that generate more human-resembling texts. If texts of different origin have high $\text{KL}_{\text{TTS}}$, it means that it is easier for a detector to separate such texts. $\text{KL}_{\text{TTS}}$ is also constrained for shorter texts, see Section 7. As for PHD, we hypothesize that generated texts of good quality should have similar PHD with human-written ones. Additionally, we compare the distributions of PHD for all datasets, see Figure 3. Again, the distributions for texts of both origin should be similar, which holds for texts from SemEval, PAN24 and MGT-1.
In the next columns, we list the statistics observed on modified texts, and for both of these the lower the better, as this reflects the similar degree of resilience of the generated and human texts to adversarial attacks. Qualitatively generated data with no bias should take values close to human.
Finally, in Table 3 we show the results of applying modern classifiers to the chosen test parts. For instance, on the datasets with low values in Table 2, a quality close to 1 can be achieved, which indicates the clear presence of detector bias towards them, or a structural feature that is too obvious for the model. It is not possible to judge the quality of the data only by achieving $F_{1}$ values close to 1, but by combining the values of the two tables, we can estimate which set has better quality data and which has lower quality data.
7 Discussion
Regarding $\text{KL}_{\text{TTS}}$, on Fig. 2 we show 4 datasets with high value of it. While GhostBuster and PAN24 received such high score due to discrepancy on texts with higher dimensions, MGTBench and DAGPap22 did it due to the difference in distributions themselves. Note also that $\text{KL}_{\text{TTS}}$ may not perform well with very short texts, since the internal method of computing PHD requires sufficiently long texts for stable computation. Therefore, we discard $\text{KL}_{\text{TTS}}$ on RuATD and AuTex23-es and TweepFake, as they do not fit the criteria, see Table 1. On top of that, it has already been shown that texts must be of sufficient length Gritsay et al. (2022) to build reliable detectors.
Analysing the values in the Table 2, we can trace the presence of sufficiently high quality data in the selected datasets. The developed attributes in aggregate are able to reflect the quality of the generated dataset from different perspectives and angles. We propose to utilise these attributes in combination with other statistical tools for evaluating data quality, e.g., Zipf’s law Powers (1998).
Presented statistics can be utilised to estimate the quality of collections and to improve them. Also, datasets that collect machine-generated content may provide utility for the two more general purposes as well. First, high-quality generated data can be utilised to evaluate the quality of the causal model during training, as one of the training objectives to improve model answers and make it more human-like. Secondly, good detectors can help to clean training sets, as large proportion of low-quality generated texts in those sets can result in emerging biases towards incorrect structure and rubbish fragments in the output of the model in the future.
8 Conclusion
In the current research, we discussed the problem of quality of datasets with AI-generated texts used for testing corresponding detectors. This problem is relevant, as the quality of test data directly influences the quality of widely used detectors. We conducted a review of datasets from competitions and scientific publications on datasets aimed at the detection of AI-generated content and proposed methods to evaluate the quality of datasets containing AI excerpts based on different structural features. We evaluated topological features, robustness to adversarial attacks, and similarity of attention patterns as estimators of data quality. We concluded that all analysed datasets fail in one another of our methods and do not allow to reliably estimate AI detectors. We encourage researchers to propose their own ways for quality assessment, which will allow to create a comprehensive system of evaluation of the detection datasets. Our work aims to contribute to a better understanding of the difference between human and machine text, which will ultimately contribute to preserving the integrity of information in the world.
9 Limitations
In our work we focused on the task of binary classification, thus suggested methods are not optimal for the task of detection of the hybrid AI-human content. Also, some methods do not work properly on short texts, however, so do the detectors.
Figure 3: PHD values on all datasets, except TweepFake and AuTex23 Spanish, texts from which were too short for proper calculation of PHD.
Appendix A Data Description
More detailed description with information on sources, topics and years of the datasets selected in this paper from competitions and research papers in Table 4
Dataset Year Themes Sources
Research papers datasets
GPT2 2019 WebText GPT-2
TweepFake 2019 Tweets Markov Chains, RNN, LSTM, GPT-2
HC3 2023 ELI5, WikiQA, Wikipedia, Medicine, Finance ChatGPT
GhostBuster 2023 Student Essays, News Articles, Creative Writing ChatGPT, Claude
MGTBench 2024 Student Essays, News Articles, Creative Writing ChatGPT. ChatGLM, Dolly, GPT4All, StableLM, Claude
MAGE 2024 Opinions, Reviews, News, QA, Story Generation, Commonsense Reasoning, Knowledge Illustration, Scientific Writing text-davinci-002, GPT-3.5, ChatGPT, LLaMA, GLM-130B, FLAN-T5, OPT, BLOOM, GPT-J-6B, GPT-NeoX-20B
M4 2024 Wikipedia, Reddit ELI5, WikiHow, PeerRead, arXiv abstract GPT-3.5, ChatGPT, Cohere, Dolly-v2, BLOOM
OutFox 2024 Student Essays ChatGPT, GPT-3.5, FLAN-T5
Shared tasks datasets
RuATD 2022 News, Social media, Wikipedia, Strategic Documents, Diaries M-BART, M2M-100, OPUS-MT, mT5, ruGPT2, ruGPT3, ruT5-Base
DAGPap 2022 Scopus papers Longformer Encoder-Decoder, GPT-3, Spinbot, GPT-Neo
AuTex 2023 Legal documents, Social media, How-to articles BLOOM, GPT-3, GPT-3.5
IberAuTex 2024 News, Reviews, Emails, Essays, Dialogues, Wikipedia, Wikihow, Tweets GPT-2, LLaMA, Mistral, Cohere, Claude, MPT, Falcon
PAN 2024 News Alpaca, BLOOM, Gemini, ChatGPT, gpt-4-turbo, LLaMA-2, Mistral, Qwen1.5, GPT-2
SemEval Mono 2024 Wikipedia, WikiHow, Reddit, arXiv, PeerRead, Student Essays GPT-3.5, GPT-4, Cohere, Dolly-v2, BLOOMz
SemEval Multi 2024 Wikipedia, WikiHow, Reddit, arXiv, PeerRead, Student Essays, News ChatGPT, GPT-3.5, GPT-4, LLaMA2, Cohere, Dolly-v2, BLOOM, Jais
MGT-1 Mono 2025 CNN, DialogSum, Wikipedia, WikiHow, Eli5, Finance, XSum, PubMed, SQuAD, IMDb, Reddit, arXiv, PeerRead text-davinci-002, GPT-3.5, ChatGPT, OPT, LLaMA3, BLOOM, FLAN-T5, Cohere, Dolly, Gemma, Mixtral
MGT-1 Multi 2025 CNN, DialogSum, Baike, WikiQA, WikiHow, Eli5, Finance, Psychology, XSum, PubMed, SQuAD, IMDb, Reddit, arXiv, PeerRead text-davinci-002, GPT-3.5, ChatGPT, gpt4o, GLM, GPT-J, GPT-Neo, OPT, LLaMA2, LLaMA3, BLOOM, FLAN-T5, Cohere, Dolly, Gemma, Mixtral, Jais
Table 4: More detailed descriptive statistics about domains and generators of the chosen datasets from competitions and research papers. ChatGPT is gpt-3.5-turbo, GPT-3.5 is text-davinci-003.
Appendix B Evaluation results of competitions
Table 5 shows the winning scores in the competitions reviewed in this paper. In the AuTex and IberAuTex competitions it was forbidden to use additional data to fine-tune the detection algorithms. In the other collections it was allowed, we can notice a high quality near perfect in them. We should note the low value of metrics on the RuATD dataset, which can be explained by the limited number of available high-quality language models in Russian during the competition.
Competition Metric Best result
RuATD Accuracy 0.820
AuTex23-en Macro-F1 0.809
AuTex23-es Macro-F1 0.708
IberAuTex Macro-F1 0.805
SemEval24 Mono Accuracy 0.975
SemEval24 Multi Accuracy 0.959
PAN24 Avg. of 5 metrics* 0.924
DAGPap22 Avg. F1-score 0.994
MGT-1 Mono Macro-F1 0.8307
MGT-1 Multi Macro-F1 0.7916
Table 5: Best results from each analysed competition. PAN24 used mean of 5 metrics, such as accuracy, F1 and other to evaluate efficiency of the system.
Figure 4: Topological Time Series on all datasets.
Appendix C Hyperparameters
Hyperparameters Values
Epochs 5*
Learning rate (LR) 5e-5
Warm-up steps 50
Weight decay 0.01
Table 6: Hyperparameters for fine-tuning mDeBERTa-base. We trained for 5 epochs with possibility of early exit.
The training was carried out on NVIDIA GeForce RTX 3090. See hyperparameters in Table 6.